OpenStudy (anonymous):

can any one tell me this question solution the fibonacci sequence is defined by 1=a1=a2 and an=an-1 + an-2 n>2. find an+1/an, for n=1,2,3,4

OpenStudy (anonymous):

you just have to grind this one out

OpenStudy (anonymous):

numbers are \[a_1=1,a_2=1,a_3=2,a_4=3\]

OpenStudy (anonymous):

\[1+\frac{1}{1}=2\] \[1+\frac{1}{1}=2\] \[2+\frac{1}{2}=\frac{5}{2}\] \[3+\frac{1}{3}=\frac{10}{3}\]

OpenStudy (anonymous):

becomes more interesting the further out in the series you go series is \[1,1,2,3,5,8,13,21,34,55,...\] where each number is the sum of the previous 2 numbers

OpenStudy (anonymous):

k.thanku

OpenStudy (anonymous):

yw